### v1

@@ 3 Scientific Discovery as Machine Learning
The problem of automating scientific discovery can be considered a machine learning problem. We cover how we can view the components of scientific discovery, as presented in Section 2, as components of machine learning techniques. There are many machine learning paradigms, so we also investigate which are the most appropriate for application to robot scientists.
Common to all machine learning techniques is that given some data to train with, we seek to learn a function that will perform in a desirable manner, and generalise beyond the given data.
In terms of the logical reasoning components of scientific discovery, forming a model ($\hat{h}$) from data is an induction problem. In machine learning settings we seek to provide good fit with our data; this could mean that we define a loss function for an optimizer, or in the case of inductive logic programming that we seek to explain positive examples and avoid inconsistencies with negative examples. To evaluate a candidate model, $\hat{h}$, we first obtain predictions from the model (deduction) and then perform some test of fit to the data, which is often a form of statistical reasoning. And finally, it is usually the case that there are many possible models that may have similar performance in the fitness tests. In which case, most machine learning techniques appeal to parsimony to choose the model which is simpler, in some defined sense. This could be the inclusion of a regularisation term in a loss function, or selecting the simplest logic program that covers the examples.
Machine learning techniques can be placed into three broad categories: supervised learning, unsupervised learning, and reinforcement learning [29]. Semi-supervised learning is sometimes also included as a fourth category, being a hybrid of supervised and unsupervised learning.
@@
  • Unsupervised learning techniques seek to find structure and extract information from unlabelled data.
@@ 3.1 Reinforcement learning is an unsatisfactory paradigm for robot scientists
We see that for pairs (A, B) we have several choices. Here we consider a few.
@@
In the first pair, our reward function could include measures of how closely the robot followed the protocol and whether it dropped equipment or created hazards such as spills. The second pair we discussed in Section 2. For third pair, we would have to construct a reward function composed of the values enumerated in Section 2. This is a daunting task to try and find a single function that incorporates each of the values. And one that goes against the advice of Kuhn who said that such a comparison would be exceedingly difficult and subjective.
@@
3.2 Supervised learning is a more appropriate paradigm
The crucial process to which we are applying machine learning is that of model improvement: given a model of an object system, how we make changes to the model such that it is more faithful to the object system. According to the scientific values presented in Section 2, there are numerous ways to evaluate this. However, in [30] “accuracy” was consistently ranked second by scientists in order of preference, only beaten by the “internal consistency” of a theory.
We argue that it is appropriate to consider scientific discovery as a supervised learning problem, both in observational and controlled experimentation. In the case of observational experiments, Medawar’s “Baconian” experiments [21], the input-output pairs will be a partition of the overall observational data. Equally, in controlled experiments, Medawar’s “Galilean experiments”, our input-output pairs will be the experimental factors and the empirical data. We can exploit various mechanisms from the field of supervised learning to obtain theories which align with the scientific values stated in Section 2, with accuracy captured in a relevant loss function. This aligns with how previous robot scientists have operated, which we cover in more detail in Section 4.
@@ 3.3 Active learning integrates agency into supervised learning
While supervised learning is the appropriate paradigm for scientific discovery, when designing robot scientists we need to integrate agency. Active learning is a specific category of supervised (or semi-supervised) learning where the learning agent chooses the next data to inform the learning. This selection policy is not the focus of the learning, and there is no requirement that the selection policy be learned. This differs from reinforcement learning, where the objective is to learn a good policy. (We could perhaps use reinforcement learning to design this agent’s policy, or alternatively use some pre-determined policy, or a combination of the two approaches.) Active learning approaches select a point in the input space, usually because the model has high uncertainty around that point. Uncertainty can arise from poor exploration of, or shallow exploitation in, the neighbourhood of that point. The agent then requests a corresponding output from an oracle. In many applications of active learning this means choosing an input-output pair from a large, extant dataset. In scientific discovery, this could be the case (querying a database of previous experimental data), or in the case of a robot scientist the oracle in question can be the experimental platform. We see that active learning shares an analogy to the scientific method, and therefore is an appropriate paradigm to choose for the design and analysis of robot scientists.
@@ 3.4 Foundation models and their use in scientific discovery
Recent developments in machine learning, driven by industrial applications and enabled by new technologies and vast amounts of data, have resulted in widespread adoption of foundation models. Foundation models are defined in [4] as
any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks.
Within the scope of this definition, most current manifestations of foundation models are large language models (LLMs)—e.g. BERT [8], GPT-3 [5]—or large multi-modal models (LMMs)—e.g. GPT-4 [23]. However, under this definition, we could also include models such as AlphaFold [14] or Evo [22], transformer-based models trained on protein structure data and genomic data respectively.
Foundation models show promise in scientific discovery applications. Besides AlphaFold there is Coscientist, a system built around GPT-4 that could autonomously design, plan and execute experiments in chemistry[3]. Coscientist exploited the general purpose nature of the foundation model to combine information from various sources and to execute code and instructions on machines to achieve its goal. In many applications there is some concern for so-called ”hallucinations” of LLMs and LMMs, i.e. claims made by the model with little or no justification or evidence. This is not a problem for a robot scientist provided the use it to generate hypotheses, as the resultant hypothesis will be tested via experiment. Hallucinations could cause problems in the experiment design phase; Coscientist mitigated the impact of hallucinations by grounding the LLM with database search, and ultimately evaluated the system’s performance using explicit criteria rather than use the LLM.
@@
These properties of LLMs and LMMs are a problem if we want to use them for “closed-loop” discovery in a robot scientist. And we also want scientific models to be interpretable and usable by other scientists, displaying the values of fruitfulness and simplicity. Foundation models often have hundreds of millions of parameters, and methods for interpreting their internal reasoning require further research before these models can be considered broadly suitable for automated discovery.
From a practical and economic perspective, LLMs and LMMs frequently require huge resources throughout the lifecycle of their development. Much focus is rightly directed to the immense electricity demands during training, but further resources are needed during research and development, data collection and storage, the construction and commissioning of hardware, and in the implementation and maintenance of LLMs and LMMs. [13].
@@ 4 Biological Systems are a Good Target for Scientific Discovery Automation
Of the scientific challenges this 21st century, understanding the biology of eukaryotic organisms ranks among the most consequential. And despite fantastic advances during the 20th century of our understanding of the fundamental components and processes in biology, and in the application of this knowledge to medicine, engineering, agriculture, etc., we are still some way off an accurate predictive model of the physiology of one organism, let alone a system of broadly applicable theories, such as those we have developed for physics.
Part of the reason why progress in biology is limited by today’s scientific methods is the diversity and complexity of the systems. Hundreds of research hours can be spent in the study of one particular gene, yet the limits of human capability and of course the economic resource available to the researcher will hamper progression to a complete understanding of the gene and its roles. Scientific discovery automation has therefore great potential in biology. And this is particularly the case when adopting the systems biology paradigm.
The cellular physiology of eukaryotes is a complex system, in the spirit of the definition given by Simon [33] that
given the properties of the parts [of the system] and the laws of their interaction, it is not a trivial matter to infer the properties of the whole.
@@
To understand complex systems it is necessary yet wholly insufficient to take a reductionist approach. Systems biology is an integrationist approach to studying biological systems. That is to say that we aim to understand how the parts and their interaction lead to the resultant behaviour of the system, be that system a cell, an organ or an entire organism.
By way of two examples of robot scientists, Adam and Eve, that were applied to the domain of biology, we will show how why this is a suitable domain for automated scientific discovery, and the importance of parsimony.
4.1 Superhuman logic and probabilistic reasoning
@@
Eve was a robot scientist designed to automate aspects of early stage drug development [34]. Eve had several modes of operation, but we focus here on the learning of quantitative structure activity relationships (QSARs). As the name suggests, they take as input the structure of a compound and predict the activity on the assay for a particular disease. Eve used a least-squares regression to learn the QSARs, which in turn were used to guide synthesis of new compounds, and further refine the QSAR. These steps are all dependent on large-scale statistical inference. Part of the reason why we need statistical inference is that the data collected from biological systems is often noisy. There is a degree of stochasticity to biological processes that are hard for humans to understand intuitively, but that machines are apt at modelling.
4.2 Applying parsimony
Adam was able to achieve autonomous discovery in part because of the parsimony of the logical theory used to generate hypotheses and evaluate them. There were many more facts and relations that could have been included in the theory, but by omitting these and building a theory that was focused on the scientific discovery task the theory was tractable and resulted only in hypotheses that were testable by the experimental apparatus available to the robot.
@@ 5 Case Study: Genesis and LGEM+
At Chalmers University in Sweden we are building a next generation robot scientist “Genesis”. Our goal is to demonstrate that the robot scientist Genesis can investigate an important area of science a thousand times more efficiently (in terms of cost and money) than human scientists.
This is an extreme challenge for AI as the number of experiments to plan and coordinate is several orders of magnitude more than the previous case studies. Achieving this goal will involve advances in automated hypothesis formation (how best to utilise background biological knowledge and models in ML, etc.), automated experiment generation (how best to optimise gain of information with cost and time constraints), laboratory robotic control, and scientific data analysis.
The scientific discovery goal of Genesis is
to develop a systems biology model of Saccharomyces cerevisiae, that is both more detailed and more accurate at predicting experimental results than any in existence.
The foundation of Genesis is a micro-fluidic system with 1000 computer-controlled micro-bioreactors (co-developed in Vanderbilt University, USA). Achieving this will be a step-change in laboratory automation as most biological labs have fewer than 10 chemostats. These micro-bioreactors are being integrated with ion-flow mass-spectroscopy (to measure metabolites at speed) and RNA-seq (to measure RNA expression levels).
To design the experiments that the robot scientist conducts, and to create and improve on the model of S. cerevisiae, we designed a modelling framework, which we present briefly now.
@@ 5.1 LGEM+: a first-order logic model
The task of scientific discovery is described by Langley [20] in generic terms that given: (a) scientific data; (b) prior knowledge about the domain; and (c) a space of candidate categories, theories, laws, or models, we seek candidates that describe or explain the data.
In Genesis’ domain, the scientific data are in the form of controlled experiments using S. cerevisiae, and empirical data arising thereof. There are many types of empirical data one could collect from such experiments. In our discovery application we focus on growth profiling and gene expression data.
Prior knowledge on S. cerevisiae is well-curated in genome-scale metabolic network models (GEMs). These are community developed models that follow a controlled vocabulary, so form a rich prior for automated scientific discovery. We chose to express the mechanisms of the biochemical pathways using first-order logic (FOL), an approach first proposed in 2001 [27]. We use a FOL structure that is grounded in the controlled vocabulary of the GEMs to express our knowledge about how entities are known to interact, for example that each reaction has reactants, products, and possible an enzyme annotation. We call this framework LGEM+; below is a brief description to illustrate the case and a more detailed explanation of the methods is in [11].
LGEM+ has five predicates: met\2, gn\1, pro\1, enz\1, and rxn\1. These are given specific semantic meaning, which is shown in Table 1. Here a cellular “compartment” refers to a component of the cellular anatomy, e.g. mitochondrion, nucleus or cytoplasm. There are seven types of clause that we included that encode the implications needed to describe phenomena such as reaction activity and gene expression. More detail on the specification of the logical theories is given in [11].
Finally, the space of candidate theories is those able to be constructed from the predicate symbols and constants relating to the cellular compartments, genes, compounds, reactions, and enzymes that could be present in yeast.
@@
[…]
@@
[…]
@@
We use automated theorem provers for first-order logic (ATPs) to perform logical inference. In comparison to previous approaches using bespoke algorithmic methods, such as MENECO [25], using an ATP removes a large part of the burden of algorithm design and simulation, particularly when it comes to abductive inference. For the reasoning tasks we use the ATP iProver [18], which was chosen due to its performance and scalability as well as completeness for first-order theorem finding. We extended iProver to include abduction inference. ATPs are designed to provide explanations, and have many tools to simplify theories to which they are applied. All these properties make ATPs a good choice for applications in scientific discovery.
@@
From the logical theory we can deduce testable facts, e.g. production of compounds. This allows us to generate input-output pairs that can be tested against truth data. The first truth data that we tested the model predictions against was single gene essentiality data for S. cerevisiae. Essential genes are those genes whose removal from the genome leads to a loss of viability for the organism. We predicted single-gene essentiality for S. cerevisiae by providing: as input (the theory $T$) the yeast genotype (including the deletion), compounds that were present in the growth medium, and compounds assumed to be ubiquitous in the cell, along with the rest of the theory containing rules for activation of reactions and formation of enzymes; and deducing the output (the goal $G$) as a binary outcome of whether every compound assessed to be essential for growth was produced.
@@
We compared these predicted pairs against empirical data for single-gene essentiality. The F1 score on the prediction task was 0.266, which was state-of-the-art for a qualitative method on that task, but still quite far away from the best quantitative models. In the case that a particular mutant is falsely predicted essential, this shows that there is room for improvement to the model, that the robot scientist needs to come up with hypotheses $H_{i}$ such that combined with the theory this hypothesis entails the goal($T\land H_{i}\vDash G$). The ATP achieves this through reverse consequence finding, using a rearrangement of the previous statement: $T\land\lnot G\vdash\lnot H_{i}$. It is also possible to steer iProver to find specific forms of H_i, though we did not do this in this case. We did filter the hypotheses found by LGEM+ to just those that might be testable via an experiment. For more details of these processes, along with further examples of deductions and abduction of hypotheses can be found in [11].
@@
The hypotheses that LGEM+ generates may well be very close together in the experimental space. The interactions between the components of such a complex system mean that a small difference in input could have a large effect on the outcome of the experiment. This is one of the benefits of using a robot scientist to execute the experiments. As we mentioned in Section 2 we can achieve higher precision with robots and also log more data that could help explain systematic errors. Yeast cultivations typically stretch over a few days, so by using robots tiredness becomes less of an issue for the human scientists involved in the project. Genesis is also equipped with a microformulator allowing for fine adjustment of input media for the cultures, to a degree that would be onerous to replicate with handheld pipettes or even a traditional liquid handling robot.

### v2

@@ 3 Scientific Discovery as Machine Learning
Satisfied that automating scientific discovery can be considered a machine learning problem, the question remains: how best should methods from the field of machine learning be applied to scientific discovery? This section covers how to view the components of scientific discovery, as presented in Section 2, as components of machine learning techniques, and investigates which of the machine learning paradigms are the most appropriate for application to robot scientists.
Common to all machine learning techniques is that given some data to train with, the goal is to learn a function that will perform in a desirable manner, and generalise beyond the given data.
In terms of the logical reasoning components of scientific discovery, forming a model ($\hat{h}$) from data is an induction problem. Machine learning seeks to form a model using information from a relevant signal to assess performance against a goal; this could mean defining a loss function for an optimiser, or in the case of inductive logic programming explaining positive examples and avoid inconsistencies with negative examples. To evaluate a candidate model, $\hat{h}$, predictions or consequences are obtained from the model (deduction) and then evaluated against the signal, often using a form of statistical reasoning. And finally, it is usually the case that there are many possible models that may have similar performance against the goal. In which case, most machine learning techniques appeal to parsimony to choose the model which is simpler, in some defined sense. This could be the inclusion of a regularisation term in a loss function, or selecting the simplest logic program that covers the examples.
Machine learning techniques can be placed into three broad categories: supervised learning, unsupervised learning, and reinforcement learning [34]; these paradigms are described below and in Figure 1. Semi-supervised learning is sometimes also included as a fourth category, being a hybrid of supervised and unsupervised learning.
@@
  • Unsupervised learning techniques seek to find structure and extract information from unlabelled data. The signal received is an intrinsic objective measure, for example data likelihood.
@@
Figure 1: Flowcharts representing high level design for: generic machine learning; supervised learning; unsupervised learning; and reinforcement learning. Unitalicised text represents inputs and outputs; italicised labels for connectors represent processes.
@@ 3.1 Reinforcement learning is an unsatisfactory paradigm for robot scientists
For pairs (A, B) there are several choices. Here we consider a few.
@@
In the first pair, the reward function could include measures of how closely the robot followed the protocol and whether it dropped equipment or created hazards such as spills. The second pair we discussed in Section 2. For third pair, we would have to construct a reward function composed of the values enumerated in Section 2. This is a daunting task to try and find a single function that incorporates each of the values. And one that goes against the advice of Kuhn who said that such a comparison would be exceedingly difficult and subjective.
@@
3.2 Supervised learning is a more useful paradigm
The crucial process machine learning is applied to in scientific discovery is that of model improvement: given a model of an object system, how to make changes to the model such that it is more faithful to the object system. According to the scientific values presented in Section 2, there are numerous ways to evaluate this. However, in Schindler [35] “accuracy” was consistently ranked second by scientists in order of preference, only beaten by the “internal consistency” of a theory.
We argue that it is useful to consider scientific discovery as a supervised learning problem, both in observational and controlled experimentation. In the case of observational experiments, Medawar’s “Baconian” experiments [26], the input-output pairs will be a partition of the overall observational data. Equally, in controlled experiments, Medawar’s “Galilean experiments”, the input-output pairs will be the experimental factors and the empirical data. When designing robot scientists, various mechanisms from the field of supervised learning can be exploited to obtain theories which align with the scientific values stated in Section 2, with accuracy captured in a relevant loss function. This aligns with how previous robot scientists have operated, which we cover in more detail in Section 4.
3.3 Semi-supervised learning
Unsupervised learning techniques, such as embedding and clustering, are used in scientific discovery applications during data representation. For example, unsupervised techniques have been used to cluster molecular dynamics data in materials science [24], and transcriptomic data in biology [17]. In each of these cases, human scientists analysed the output of the unsupervised learning to draw conclusions.
As discussed above, a supervised learning design allows robot scientists to close the discovery loop and not require human scientists to analyse results. However, semi-supervised learning—integrating unsupervised techniques into supervised learning—allows for contextual information and the structure of theories to be exploited in the discovery task, as well as addressing the issue that the experimental space is sparsely annotated.
Semi-supervised techniques can exploit the background knowledge that is available to the robot scientist to make better predictions and learn more efficiently. Gleaves et al. [13] used a semi-supervised framework to improve synthesisability models in materials science, giving evidence that a semi-supervised approach to learning could provide the most promise for scientific discovery applications.
@@ 3.4 Active learning integrates agency into supervised learning
While supervised learning is the appropriate paradigm for scientific discovery, the design of a robot scientist must integrate agency. Active learning is a specific category of supervised learning where the learning agent chooses the next data point (input) for which a label (output) has not been observed. This selection policy is not the focus of the learning, and there is no requirement that the selection policy be learned. This differs from reinforcement learning, where the objective is to learn a good policy. (Reinforcement learning could be used to design this agent’s policy, or alternatively use some pre-determined policy, or a combination of the two approaches.)
Active learning approaches select a point in the input space that has not been observed. Points are selected, usually either because the model has high uncertainty around that point, or to increase the diversity of the dataset. Uncertainty can arise from poor exploration of, or shallow exploitation in, the neighbourhood of that point. The agent then requests a corresponding output. In many applications of active learning this means asking for a human or expert annotation. In scientific discovery, this could be the case, or in the case of a robot scientist it can use its experimental platform to perform an experiment to get output data. We see that active learning shares an analogy to the scientific method, and therefore is an appropriate and useful paradigm to choose for the design and analysis of robot scientists. Both Adam and Eve used active learning by searching the hypothesis space and executing experiments to then improve the scientific model. This is covered in more detail in Section 4.1.
@@ 3.5 Foundation models and their use in scientific discovery
Recent developments in machine learning, driven by industrial applications and enabled by new technologies and vast amounts of data, have resulted in widespread adoption of foundation models. Foundation models are defined in [5] as:
“any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks.”
Within the scope of this definition, most current manifestations of foundation models are large language models (LLMs)—e.g. BERT [10], GPT-3 [6]—or large multi-modal models (LMMs)—e.g. GPT-4 [28]. However, this definition also encompasses models such as Evo [27] or ESM3 [15], transformer-based models trained on genomic data and protein data respectively.
Foundation models show promise in scientific discovery applications. Besides AlphaFold there is Coscientist, a system built around GPT-4 that could autonomously design, plan and execute experiments in chemistry [4]. Coscientist exploited the general purpose nature of the foundation model to combine information from various sources and to execute code and instructions on machines to achieve its goal. In many applications there is some concern for so-called “hallucinations” of LLMs and LMMs, i.e. claims made by the model with little or no justification or evidence. This is not a problem for a robot scientist provided they use it to generate hypotheses, as the resultant hypothesis will be tested via experiment. Hallucinations could cause problems in the experiment design phase; Coscientist mitigated the impact of hallucinations by grounding the LLM with database search, and ultimately evaluated the system’s performance using explicit criteria rather than use the LLM.
@@
These properties of LLMs and LMMs are a problem for use in “closed-loop” discovery in a robot scientist. Scientific models should be interpretable and usable by other scientists, displaying the values of fruitfulness and simplicity. Foundation models often have hundreds of millions of parameters, and methods for interpreting their internal reasoning require further research before these models can be considered broadly suitable for automated discovery.
From a practical and economic perspective, LLMs and LMMs frequently require huge resources throughout the life cycle of their development. Much focus is rightly directed to the immense electricity demands during training, but further resources are needed during research and development, data collection and storage, the construction and commissioning of hardware, and in the implementation and maintenance of LLMs and LMMs [18].
@@ 4 Biological Systems are a Good Target for Scientific Discovery Automation
Of the scientific challenges in this 21st century, understanding the biology of eukaryotic organisms ranks among the most consequential. Despite fantastic advances during the 20th century of our understanding of the fundamental components and processes in biology, and in the application of this knowledge to medicine, engineering, agriculture, etc., we are still some way off an accurate predictive model of the physiology of one organism, let alone a system of broadly applicable theories, such as those developed for physics.
Part of the reason why progress in biology is limited by today’s scientific methods is the diversity and complexity of the systems. Hundreds of research hours can be spent in the study of one particular gene, yet the limits of human capability and of course the economic resource available to the researcher will hamper progression to a complete understanding of the gene and its roles. Scientific discovery automation has therefore great potential in biology. This is particularly the case when adopting the systems biology paradigm.
The cellular physiology of eukaryotes is a complex system, in the spirit of the definition given by Simon [38] that:
“given the properties of the parts [of the system] and the laws of their interaction, it is not a trivial matter to infer the properties of the whole.”
@@
To understand complex systems it is necessary, yet wholly insufficient, to take a reductionist approach. Systems biology is an integrationist approach to studying biological systems. The aim is to understand how the parts and their interaction lead to the resultant behaviour of the system, be that system a cell, an organ or an entire organism.
By way of two examples of robot scientists, Adam and Eve, that were applied to the domain of biology, we will show why this is a suitable domain for automated scientific discovery.
4.1 Discovery through superhuman logic and probabilistic reasoning
@@
Eve was a robot scientist designed to automate aspects of early stage drug development [40]. Eve had several modes of operation, but we focus here on the learning of quantitative structure activity relationships (QSARs). As the name suggests, they take as input the structure of a compound and predict the activity on the assay for a particular disease. Eve used a least-squares regression to learn the QSARs, which in turn were used to guide synthesis of new compounds, and further refine the QSAR. These steps are all dependent on large-scale statistical inference, partly because data collected from biological systems are often noisy. There is a degree of stochasticity to biological processes that are hard for humans to understand intuitively, but that machines are apt at modelling.
@@ 5 Case Study: Genesis and LGEM+
At Chalmers University in Sweden we are building a next-generation robot scientist “Genesis”. Our goal is to demonstrate that the robot scientist Genesis can investigate an important area of science a thousand times more efficiently (in terms of cost and money) than human scientists.
This is an extreme challenge for AI as the number of experiments to plan and coordinate is several orders of magnitude more than the previous case studies. Achieving this goal will involve advances in automated hypothesis formation (how best to utilise background biological knowledge and models in ML, etc.), automated experiment design (how best to optimise gain of information with cost and time constraints), laboratory robotic control, and scientific data analysis.
The scientific discovery goal of Genesis is to develop a systems biology model of S. cerevisiae, that is both more detailed and more accurate at predicting experimental results than any in existence. An outline of Genesis’ discovery process is shown in Figure 2.
Figure 2: Flowcharts representing the scientific discovery process of Genesis. The robot scientist starts with a systems biology theory, contructed from community knowledge. After using a computational model (e.g. LGEM+) to form hypotheses, Genesis will design and run lab experiments using its hardware, and will take measurements of phenotype using automated procedures, e.g. for metabolomics. Simulated phenotype will be compared against the observed phenotype to generate information used to refine the theory, and the cycle will begin again with the improved theory.
The foundation of Genesis is a micro-fluidic system with 1000 computer-controlled micro-bioreactors (or chemostats) co-developed in Vanderbilt University. Achieving this will be a step-change in laboratory automation as most biological labs have fewer than 10 chemostats. These micro-bioreactors are being integrated with ion-flow mass-spectroscopy (to measure metabolites at speed) and RNA-seq (to measure RNA expression levels).
To design the experiments that the robot scientist conducts, and to create and improve on the model of S. cerevisiae, we designed a modelling framework, which we present briefly next.
@@ 5.1 LGEM+: a first-order logic model
The task of scientific discovery is described by Langley [25] in generic terms that given: (a) scientific data; (b) prior knowledge about the domain; and (c) a space of candidate categories, theories, laws, or models, scientists seek the candidates that describe or explain the data.
In Genesis’ domain, the scientific data are in the form of controlled experiments using S. cerevisiae and resultant empirical data. There are many types of empirical data one could collect from such experiments. In our discovery application we focus on metabolomics and gene expression data.
Prior knowledge on S. cerevisiae is well-curated in genome-scale metabolic network models (GEMs). These are community developed models that follow a controlled vocabulary, so form a rich prior for automated scientific discovery. We chose to express the mechanisms of the biochemical pathways using first-order logic (FOL), an approach first proposed in 2001 [32]. We use a FOL structure that is grounded in the controlled vocabulary of the GEMs to express knowledge about how entities are known to interact, for example that each reaction has reactants, products, and possibly an enzyme annotation. We call this framework LGEM+; below is a brief description to illustrate the case and a more detailed explanation of the methods is in [14].
LGEM+ has five predicates: met/2, gn/1, pro/1, enz/1, and rxn/1. These are given specific semantic meaning, which is shown in Table 1. Here a cellular “compartment” refers to a component of the cellular anatomy, e.g. mitochondrion, nucleus or cytoplasm. There are seven types of clause that we included that encode the implications needed to describe phenomena such as reaction activity and gene expression. More detail on the specification of the logical theories is given in [14].
Finally, the space of candidate theories is those able to be constructed from the predicate symbols and constants relating to the cellular compartments, genes, metabolites, reactions, and enzymes that could be present in yeast.
@@
gn/1 gene identifier “Gene X is expressed.”
pro/1 protein complex identifier “Protein complex X is available (in every
@@
enz/1 enzyme category identifier “Enzyme category X is available.”
rxn/1 reaction “There is positive flux through reaction X.”
@@
We use automated theorem provers for first-order logic (ATPs) to perform logical inference. ATPs are software that can automatically prove or disprove logical statements, by applying logical inference to a set of axioms. In comparison to previous approaches using bespoke algorithmic methods, such as MENECO [30], using an ATP removes a large part of the burden of algorithm design and simulation, particularly when it comes to abductive inference. For the reasoning tasks we use the ATP iProver [22], which was chosen due to its performance and scalability as well as completeness for first-order theorem finding. We extended iProver to include abduction inference. ATPs are designed to provide explanations, and have many tools to simplify theories to which they are applied. All these properties make ATPs a good choice for applications in scientific discovery.
@@
From the logical theory, the ATP can deduce testable facts, e.g. production of metabolites. This allows us to generate input-output pairs that can be tested against truth data. The first truth data the model predictions were tested against were single gene essentiality data for S. cerevisiae. Essential genes are those genes whose removal from the genome leads to a loss of viability for the organism. Single-gene essentiality was predicted for S. cerevisiae by providing: as input (the theory $T$) the yeast genotype (including the deletion), metabolites that were present in the growth medium, and metabolites assumed to be ubiquitous in the cell, along with the rest of the theory containing rules for activation of reactions and formation of enzymes; and deducing the output (the goal $G$) as a binary outcome of whether every metabolite assessed to be essential for growth was produced.
@@
Predictions were compared against empirical data for single-gene essentiality. The F1 score on the prediction task was 0.266, which was state-of-the-art for a qualitative method on that task, but still quite far away from the best quantitative models. In the case that a particular mutant is falsely predicted essential, this shows that there is room for improvement to the model, that the robot scientist needs to come up with hypotheses $H_{i}$ such that combined with the theory this hypothesis entails the goal ($T\land H_{i}\vDash G$). The ATP achieves this through reverse consequence finding, using a rearrangement of the previous statement: $T\land\lnot G\vdash\lnot H_{i}$. It is also possible to steer iProver to find specific forms of $H_{i}$, though we did not do this in this case. We filter the hypotheses found by LGEM+ to just those that might be testable via an experiment, and used flux balance analysis to select plausible candidates. Hypotheses that correct errors with minimal effect on the current model (produce as few as possible new metabolites) are selected. This is in accordance with the principle of parsimony as discussed in Section 2.1.3. For more details of these processes, along with further examples of deductions and abduction of hypotheses can be found in [14].
@@
The hypotheses that LGEM+ generates may well be very close together in the experimental space. The interactions between the components of such a complex system mean that a small difference in input could have a large effect on the outcome of the experiment. This is one of the benefits of using a robot scientist to execute the experiments. As we mentioned in Section 2 we can achieve higher precision with robots and also log more data that could help explain systematic errors. Yeast cultivations typically stretch over a few days, so by using robots, tiredness becomes less of an issue for the human scientists involved in the project. Genesis is also equipped with a microformulator allowing for fine adjustment of input media for the cultures, to a degree that would be onerous to replicate with handheld pipettes or even a traditional liquid handling robot.